#4 The probability of dinning in Restaurant A is 0.55 while the probability of dinning in restaurant B is 0.35.

The probability of dinning in both restaurants is 0.20. What is the probability of dinning in either restaurant A or B
Answer #1

P(A) = 0.55, P(B) = 0.35, P(both) = 0.2

P(Either) = P(A) + P(B) - P(both) = 0.55 + 0.35 - 0.2 = 0.7

Hence, 0.7 or 70% is the probability of dinning in either restaurant A or B.

Please comment if any doubt. Thank you.